By what percent the energy of satallite has to be increased to shift it from an orbit of radius R to 3/2R
Ans 33.3%
1
Expert's answer
2019-04-05T10:27:44-0400

The energy of satellite in the first case:


"E_1=\frac{GMm}{2R}"

In the second case:


"E_2=\frac{GMm}{2(1.5R)}=\frac{GMm}{3R}"

"\Delta E=-\frac{GMm}{3R}-(-\frac{GMm}{2R})=\frac{GMm}{6R}"

% increase:


"\frac{\frac{GMm}{6R}}{\frac{GMm}{2R}} 100 \%=33.3 \%"